# GET: Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v1

@@ Appendix 0.E Visualization of Different Pre-trained Models
It can be seen that different backbone models focus on different areas and they lack attention to more detailed and discriminative regions, making it challenging to distinguish visually similar categories. We argue that this is due to the model being trained solely on visual information from the dataset, making it challenging to generalize well to visually similar classes. Though visually similar, their text information tends to be different, promoting us to introduce text information into the task. In the meanwhile, the potential of current GCD methods heavily relies on the generalization ability of pre-trained models, prompting us to select a more robust and realistic pre-training model. As a large-scale model, CLIP shows strong generalization ability on downstream tasks and strong multi-modal potential due to its image-text contrastive training. Thus introducing CLIP into the GCD task not only enhances the basic class generalization ability but also builds a bridge to use use text information.
Appendix 0.F Qualitative Results
0.F.1 Cluster results of GET
As shown in Fig. 11, we present the comparative cluster accuracy between our multi-modal approach and previous single-modal methods on some visually similar classes in CUB datasets. It is worth noting that relying solely on visual information, even with a powerful CLIP backbone, the previous method (SimGCD-CLIP) still struggles to differentiate some categories, resulting in empty clusters. However, leveraging the rich and discriminative text information of categories, our GET achieves more accurate classification results on CUB without any empty clusters across all categories, demonstrating the importance of multi-modal information in the GCD task. Furthermore, we showcase the clustering results of SimGCD-CLIP (see in Fig. 14) and our GET  (see in Fig. 15) for the 170th class, “Mourning Warbler”, in the CUB dataset. SimGCD-CLIP relies solely on visual information to categorize birds based on shape and posture, the model categorizes many visually similar samples as “Mourning Warbler”. Our approach, by incorporating text information, enhances the model’s discriminative ability and correctly identifies all instances of the “Mourning Warbler” class, achieving 100% classification accuracy for this visually challenging category.
Figure 11: Cluster accuracy of SimGCD-CLIP and our GET on some visually similar classes in CUB datasets. GCD methods relying solely on a single visual modality result in empty clusters(highlighted by red boxes); Our multi-modal approach GET avoids empty clusters and achieves higher classification accuracy.
0.F.2 Attention map visualization of GET
@@
We visualize the attention map of class tokens in Fig. 12. Compared to SimGCD-CLIP, our approach additionally focuses on the feather texture of birds, which is crucial for distinguishing visually similar fine-grained bird species. With the assistance of text information, the attention maps of our visual branch become more refined, focusing on more discriminative regions.
@@
0.F.3 The t-SNE visualization
We randomly sample 20 classes from the 200 classes in the CUB dataset and perform the t-SNE projections of visual and text features. As shown in Fig. 13, both visual and text features of our method exhibit clear and compacter clusters.
@@ 0.F.3 The t-SNE visualization
Appendix 0.G Broader Impact and Limitations
Our approach introduces text information into the GCD task through the text embedding synthesizer module, extending the GCD to a multi-modal paradigm. Our method achieves state-of-the-art performance on all benchmark datasets and demonstrates the generalizability of TES in both parameterized and non-parameterized GCD methods. Our dual-branch multi-modal joint training framework with a cross-modal instance consistency objective provides an effective approach for leveraging multimodal information. A limitation of our approach is that we treat visual and text information as equally important. In fact, some samples may have richer and more discriminative visual information than textual information, and vice versa. A more appropriate approach might involve enabling the model to adaptively leverage multimodal information, autonomously assessing which modality’s information is more crucial. We will delve deeper into this aspect in our future work.
Figure 14: SimGCD-CLIP cluster results visualization for class “Mourning Warbler” in CUB dataset. SimGCD-CLIP categorizes birds based on shape and posture and incorrectly identifies many visually similar categories, resulting in a clustering accuracy of 6.7% for class “Mourning Warbler”.
Figure 15: Cluster results visualization for class “Mourning Warbler” in CUB dataset of our GET. Our method uses multi-model information, achieving 100% classification accuracy for this visually challenging category.

### v2

@@ 5.5 Different pre-trained models
A key observation is that though promising results have been achieved, different backbones, even powerful CLIP, still perform inferiorly on distinguishing certain visually similar classes, such as the classes in all fine-grained datasets. We argue that this is due to current methods only utilize a single visual modality of information, another modality may potentially compensate for the lack of discriminative ability. In the meanwhile, the potential of current GCD methods heavily relies on the generalization ability of pre-trained models, prompting us to select a more robust and realistic pre-training model. As a large-scale model, CLIP shows strong generalization ability on downstream tasks and strong multi-modal potential due to its image-text contrastive training, thus we decide to introduce it into the GCD task. This not only unleashes the latent potential performance of existing methods but also serves as a bridge for us to leverage multi-modal information.
5.6 Qualitative results
@@
Attention map visualization of GET.  We visualize the attention map of class tokens in Fig. 8. Compared to SimGCD-CLIP, our approach additionally focuses on the feather texture of birds, which is crucial for distinguishing visually similar fine-grained bird species. With the assistance of text information, the attention maps of our visual branch become more refined, focusing on more discriminative regions.
@@
The t-SNE visualization.  We randomly sample 20 classes from the 200 classes in the CUB dataset and perform the t-SNE projections of visual and text features. As shown in Fig. 9, the features of SimGCD-CLIP are not well separated, while both visual and text features of our method exhibit clear and compacter clusters. We also provide the cluster results of GET in the Appendix.
@@
6 Conclusions
In this work, we propose to leverage multi-modal information to solve the GCD task. In particular, we introduce a text embedding synthesizer to generate pseudo text embeddings for unlabelled data. Our text embedding synthesizer module makes it possible to use CLIP’s text encoder, thus unlocking the multi-modal potential for the GCD task. Meanwhile, we use a dual-branch training strategy with a cross-modal instance consistency objective, which facilitates collaborative action and mutual learning between the different modalities, thus generating more discriminative classification prototypes. The superior performance on multiple benchmarks demonstrates the effectiveness of our method.
Limitations and future works.  A limitation of our approach is that we treat visual and text information as equally important. In fact, some samples may have richer and more discriminative visual information than textual information, and vice versa. A more appropriate approach might involve enabling the model to adaptively leverage multimodal information, autonomously assessing which modality’s information is more crucial. We will delve deeper into this aspect in our future work.
Broader impact.  Our approach introduces text information into the GCD task through a novel text embedding synthesizer module, extending the GCD to a multi-modal paradigm. We believe that the introduction of TES will encourage future research in solving GCD in a multi-modal manner.
7 Acknowledgments